23. A game consists of tossing a one rupee coin 3 times and noting its outcome each time. Hanif wins if all the tosses give the same result i.e., three heads or three tails, and loses otherwise. Calculate the probability that Hanif will lose the game.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
The total number of outcomes = 8 (HHH, HHT, HTH, THH, TTH, HTT, THT, TTT) Total outcomes in which Hanif will lose the game = 6 (HHT, HTH, THH, TTH, HTT, THT) P (losing the game) = 6/8 = ¾ = 0.75